# Towards Adapting Federated & Quantum Machine Learning for Network Intrusion Detection: A Survey

### v1

@@ IV-I1 DDoS (Distributed Denial of Service) Attacks
A DDoS attack disrupts the services provided by the systems by flooding them with excessive requests, rendering them unable to function properly. A DDoS attack can have a severe impact on the federated learning process, slowing the model aggregation process or exhausting the limited resources of edge devices [93]. Alhasawi and Alghamdi in [94] presented a novel solution, Federated Learning for Decentralized DDoS Attack Detection (FL-DAD) using CNN, to detect DDoS attacks in IoT networks. They also mentioned the challenges while implementing their solutions which include achieving consistent accuracy due to heterogeneous data, resource constraints on edge devices, communication overhead, anomalous data intricacies and computation complexities. In spite of these challenges, FL-DAD constantly achieved an accuracy of 98% for various DDoS classes, outperforming traditional centralized methods.
Similarly, Fotse et al. in [93] developed FedLAD, to detect DDoS attacks in large-scale software-defined networks (SDN). XGBoost classifier was implemented to detect DDoS attacks within the SDN. They implemented three different aggregation strategies and compared the results for three different datasets. The results showed that Astraes and Ranking Client outperform FedAvg in most performance metrics and the results were almost similar to the centralized approach. This method helps in minimal resource consumption by distributing the detection process to multiple controllers enhancing detection speed. A few limitations of FedLAD are model accuracy being dependent on aggregation techniques and the necessity to protect the model against data poisoning threats.
@@ IV-I2 MITM (Man in the Middle) Attacks
MITM attacks are responsible for data theft, unauthorized access and system integrity damage. Various techniques such as ARP spoofing, session hijacking, DNS spoofing and SSL stripping are used for MITM attacks [95]. Satpathy et al. in [96] conducted a performance analysis using FL models on Gradient Boosting Machines (FL-GBM) and Long Short-Term Memory (FL-LSTM) combining them with PCA (Principle Component Analysis) for MITM attacks detection. The results showed that FL-GBM outperformed FL-LSTM in terms of accuracy and F1 score.
Similarly, Pathmendre et al. in [97] presented a novel security framework CodeNexa for MITM attack detection. Existing security mechanisms include hash functions, digital signatures and watermarking to protect the integrity of the data. These traditional methods though effective have several limitations. Hash functions verify static data integrity but are not able to adapt to the dynamic and evolving nature of FL [98]. Digital signatures maintain the authenticity of the data but the attacker can intercept and manipulate both data and signature during transmission [99]. Similarly, watermarks ensure ownership verification but fail to protect continuous model integrity [100]. The proposed method does not include cryptography; instead, it enhances model integrity by dynamic metric verification such as accuracy, precision, recall and area under the curve (AUC). During global model aggregation, the stored metrics are compared with newly calculated metrics from local updates before aggregating the model. The model update is accepted only if a certain threshold is met. The metrics need to be transmitted to the global server along with model weights which remains a challenge as it increases the communication overhead.
@@ IV-I3 Botnet Attacks
Alhasawi et al. [94] Horizontal Weighted FedAvg Cross-device CICIDS2017 CNN DDoS
Fotse et al. [93] Horizontal Astraes, Ranking Client, FedAVG Cross-silo InSDN, CICDDoS2019, CICDoS2017 XGBoost DDoS
Satpathy et al. [96] Horizontal FedAvg Cross-device UNSW-NB15 GBM, LSTM with PCA MITM
Pathmendre et al. [97] Horizontal CodeNexa (validation-based) Cross-device MNIST CNN MITM
@@ V Quantum Machine Learning (QML)
Quantum computing is rapidly evolving and getting popular as it leverages the principles of quantum mechanics (i.e., superposition, entanglement, decoherence, and quantum interference) that process information in a different way as compared to classical computers [104]. Unlike classical bits (i.e., 0 or 1), quantum bits (qubits) can exist in a superposition of both the states which helps in parallel computation across a vast solution space. A qubit can be in the state $|\psi\rangle=|0\rangle$ or $|\psi\rangle=|1\rangle$, and it can also be in a superposition state, i.e.,
@@
The extreme values of the qubits are 0 and 1, but they can be anything in between, representing a state formed by the combination of 0 and 1. This is a superposition of two values. A qubit is capable of representing a larger amount of data since every state of a qubit can be mapped to different pieces of data. This makes a single qubit solve more computations as compared to many regular classical bits that work simultaneously. Quantum computers rely on quantum Turing machines for computation, in contrast to classical Turing machines in classical computers [105]. The runtime of an algorithm is measured in terms of the number of elementary operations $N$ for both classical and quantum computation [106]. In classical computing, the same problem can belong to NP but not P, meaning it can be verified in polynomial time but has no known polynomial-time solving algorithm on a classical machine, while it can belong to class BQP in quantum machines when defined on Hilbert spaces [107]. These distinctive features of quantum computing enable them to solve complex optimization and machine-learning problems involving high-dimensional data and intricate patterns.
Quantum computing relies on Dirac notation (also called bra-ket notation) to efficiently describe and manipulate quantum states [108]. A quantum state is written as ket, denoted by $|\psi\rangle$. This represents a column vector in a complex vector space (Hilbert Space). The corresponding bra, denoted by $\langle\phi|$, is a conjugate transpose of a ket. It represents a row vector and is used to form inner products and matrix operations.
@@ V-A Quantum Gates and Circuits
However, quantum gates operate on qubits that can exist in a superposition of 0 and 1. Quantum gates are reversible and linear. The output of these gates gives the probability of qubits being 0 or 1. The gates are classified based on the number of qubits they operate on. Gates that operate on a single qubit are called single-qubit gates, those on two qubits are called two-qubit gates, and those on three qubits are called three-qubit gates [109]. The most known single-qubit gates are Pauli X gate, Pauli Z gate and Hadamard (H) gate:
@@
A quantum circuit is created by combining these gates to get desired outputs. The example of bell state [110] shows how quantum gates can be used to create a quantum circuit and how quantum circuits can be used to perform various operations on quantum bits.
Traditional quantum circuits are static - they are designed to perform fixed computations [111]. The concept of parameterized quantum circuits (PQCs), is introduced where certain gates in the circuit contain learnable parameters, similar to weights in classical neural networks. The PQCs are also called variational circuits. In QML, a PQC acts like a model, analogous to a neural network. Classical data are encoded into quantum states using encoding circuits or quantum feature maps [112]. The variational circuit applies transformations based on trainable parameters on these quantum-encoded data. After measurement, classical values are used to calculate a loss function. Various classical optimizers (like gradient descent, Adam) can be used to adjust the parameters based on the loss function. The rotation angles and parameters in the quantum gates in a variational circuit are learned during QML. This learning occurs by minimizing the loss function, similar to classical settings.
@@ V-B Feature Encoding for Quantum Intrusion Detection
The foremost step before implementing QML techniques is to map classical features into quantum states so that QML algorithms can operate on those features. Feature mapping is one of the most crucial steps in QML. Network traffic data used in IDS is high-dimensional, noisy and non-linearly correlated. The quantum feature maps exploit exponentially high-dimensional Hilbert space to encode classical features into quantum states [113], which allow quantum classifiers or kernels to detect subtle and complex patterns indicative of intrusions or attacks.
@@
The common feature mapping techniques include [114]:
@@ V-C QML in Intrusion Detection
With growing interconnectedness between devices [115], the growth in data has been exponential. In the modern world, we rely on machine learning to help systems make automatic and smart decisions. Machine learning algorithms have made the process of decision-making very fast. One of the most common examples is detecting spam in emails. It could take days for a human to filter spam from a collection of emails by manually checking them whereas a ML algorithm can filter these spam in seconds. But as the amount and complexity of data is constantly growing, classical ML sometimes struggles.
@@
To address this question, it is important to understand the limitations of classical ML algorithms. Classical ML models work by processing combinations of input at a time which makes the model learn patterns point by point [116]. This becomes computationally heavy if the data is huge and patterns are complex.
@@
These principles led to the integration of quantum algorithms into traditional machine learning workflows [117]. QML models such as quantum support vector machines (QSVMs) [118], variational quantum circuits (VQCs) [119], and quantum kernel methods [120], can provide exponential or polynomial speed-ups in processing data. In the field of cyber security, QML has been investigated for anomaly detection, traffic pattern analysis and malware detection, where enhanced feature extraction and generalization capabilities are critical [104, 121, 122].
As technologies in quantum computing evolve, researchers are exploring the integration of QML into IDS. Most QML-based IDS fall into two main categories - VQC-based and quantum kernel-based [112]. VQC consists of a PQC that uses a sequence of quantum gates controlled by tunable classical parameters. The parameters of PQC are optimized in classical computers to reduce the cost function. On the other hand, kernel-based methods make use of quantum computers to encode the data into high-dimensional Hilbert spaces. The similarities between data points are measured using inner products of quantum states. The quantum kernel captures how similar the two inputs are in this space. A decision boundary is defined by suitable quantum measurement, allowing complex patterns (like anomalies) to be separated linearly in the transformed space.
@@ V-C1 Supervised Quantum Algorithm
  • Quantum Support Vector Machine (QSVM): QSVM is the counterpart of classical SVM. SVM is based on a hyperplane that separates data points of different classes with maximum margin. QSVM is based on a similar approach, but it performs the computation in a high-dimensional Hilbert space using quantum states and quantum kernels. Quantum computers can evaluate the inner products faster, leading to faster evaluation of kernel matrix [123]. The QSVM can encode classical data into exponentially large Hilbert spaces [113], enabling the separation of classes that are not linearly separable in classical classes. The attacker will make sure that the anomalous data looks like normal data to not get caught, which makes classical SVM separate the normal/anomalous class challenging. However, in QSVM, the encoded data in high-dimensional space makes a clear boundary between normal/anomalous data facilitating more accurate and efficient classification.
@@
  • Quantum Neural Network (QNN): A QNN is a variational circuit with learnable parameters, which are to be optimized while training [124]. The gates in variational circuits are controlled by tunable parameters. These tunable parameters are adjusted during training like weights in classical ML. Fig. 8 shows an overview of QML in intrusion detection systems.Designing quantum circuits for intrusion detection include several key steps:
@@
  • Quantum k-Nearest Neighbor (QkNN): The classical kNN algorithm is based on a distance metric (e.g. Euclidean distance). This algorithm identifies the k closest points (neighbors) to a query data point based on a distance metric and uses their label to predict the label of the query data. However, when the data size and dimensionality increase, the classical kNN becomes computationally very expensive due to the need to calculate distances between all the data and a query data point. In QkNN, all training data can be encoded into a superposition of quantum states. Unlike in classical kNN where looping is required on $n$ states one by one, the quantum algorithm can process all of them at once in a single operation using quantum parallelism. The distance in QkNN can be measured using fidelity [125], which is related to inner products and is given by:
@@
    The fidelity is computed using quantum circuits such as swap tests. Mathematically, fidelity quantifies the overlap between two quantum states, with values ranging from 0 (completely different) to 1 (identical). This measure is crucial for distinguishing subtle differences in network behavior, especially when anomalous traffic closely resembles normal traffic patterns to avoid detection. QkNN can alos be implemented using various distance metrics (such as Hamming and Euclidean distance) [126, 127]. In IDS, the query network traffic can be compared with the labeled traffic data using fidelity or various other distance metrics to classify it as normal or anomalous. QkNN can help detect anomalous network data faster as compared to classical kNN due to quantum parallelism.
  • Hybrid Quantum Classical Neural Network (HQCNN): The HQCNN is a combination of quantum and classical neural networks [128, 129]. In HQCNN, the feature dimension of classical data is reduced by passing through a classical neural network. The output of the first classical neural network layer acts as input to the quantum circuit. The output is encoded as quantum states using methods like angle or amplitude encoding so that quantum algorithms can operate on it. The quantum operation happens in the second layer. The output of the quantum layer is measured and fed back to the classical neural network layer to get the desired output. This method can efficiently detect anomalies in network traffic where network features have interdependent patterns. Quantum entanglement can help capture dependencies between different types of network behaviors.
@@ V-C2 Unsupervised Quantum Algorithm
  • Quantum Autoencoder (QAE): Classical autoencoders encode classical data into a lower dimension and learn to reconstruct the original classical data from this lower dimension [130]. The training is done between the input and output of this algorithm to reduce the reconstruction loss. Similarly, QAE encodes the quantum data stored on n qubits into a smaller quantum register of $m<n$ qubits [131](see Fig. 9). In quantum computing, the qubits cannot be simply erased/discarded due to quantum operations being unitary (i.e., reversible). Suppose an initial quantum state living in Hilbert space is given by:
@@
    In [132], Aïmeur et al. introduced thee Grover-based quantum subroutines to accelerate classical clustering tasks. These subroutines include quant-find-max (used to find the pair of points in the dataset with maximum distance), quant-find-c-smallest-values (which identifies the $c$ nearest neighbors of a given point), and quant-cluster-median (a novel quantum routine that determines the median of a point set by computing the total distance from each point to all others and then applying Grover’s minimum search to identify the point with the smallest overall distance). Using these primitives, authors developed three clustering algorithms: quantum divisive clustering, $k$-medians and construction of a $c$-neighborhood graph.
@@
    All these algorithms are based on Grover’s algorithm [133] which attains quadratic speedup compared to classical methods. Suppose there are $N$ unsorted items, classically it would take $N/2$ steps to search items on average and $N$ steps in the worst case. Grove’s algorithm can find the solution in $\sqrt{N}$ steps by leveraging quantum parallelism. The ability of quantum computing to handle high-dimensional data and quadratic speedup achieved in Grover’s algorithm shows its potential in network attack detection, particularly suitable for real-time and adaptive security applications.
@@
  • Quantum Ensemble Methods: They extend classical ensemble methods, such as bagging, boosting, and stacking, into the quantum domain [134]. In classical learning, many weak learners are used to develop a strong learner that outperforms a single model. Similarly, in quantum computing, quantum ensemble techniques leverage the quantum probabilistic nature and parallelism of quantum computing to produce diverse models and aggregate them to develop a strong model with enhanced performance [135]. Quantum ensemble methods can capture richer patterns in network traffic using fewer resources compared to the classical model. Classical ensemble methods often struggle with high-dimensional and complex network flow data whereas quantum models can process high-dimensional and complex data in exponentially smaller space which makes quantum ensemble methods more suitable for IDS. It can help reduce false positives and accurately detect anomalies in the network.
@@
  • Quantum Boltzmann Machine (QBM): It is a quantum-enhanced version of the classical Boltzmann machine that leverages the principles of quantum mechanics. The classical energy function is replaced with a quantum Hamiltonian, and the system’s state is described by a thermal quantum density matrix in a QBM [22]. This allows a QBM to model high dimensional probability distribution more efficiently as compared to the classical Boltzmann machine. The patterns in network traffic are very intricate. These patterns can make classical Boltzmann machine hard to learn complex patterns. The QBM can help reduce false positives and detect real-time anomalies due to high-dimensional feature learning which can be very advantageous in IDS.
@@
  • Quantum Generative Adversarial Network (qGAN): The qGAN combines the power of GANs learning paradigm and quantum computation [124]. The qGANs use a quantum generator $G_{\Theta}$ to generate synthetic data distributions that closely resemble original data and discriminator $D_{\phi}$ to distinguish between original and generated data distributions. Both generator and discriminator are trained together in a min-max game where $G_{\Theta}$ goal is to fool the discriminator by generating so realistic data that $D_{\phi}$ cannot tell it’s fake and $D_{\phi}$’s goal is to classify the data as real or fake. Lloyd et al. in [136] first proposed the concept of qGAN, showing that quantum generators could be more expressive than classical ones, especially when operating in high-dimensional Hilbert spaces. This adversarial learning can be particularly advantageous for intrusion detection, where labeled data is scarce and anomalies are rare, making it suitable for unsupervised or semi-supervised learning. By learning the distribution of normal traffic, qGANs can flag any deviations from normal behavior as a potential threat.
@@
  • Quantum PCA (QPCA): QPCA is a quantum analogue of classical PCA, which is designed to reduce the dimensionality of quantum data while preserving the most significant features. It was first introduced by Lloyd et al. [137], who proposed an efficient quantum algorithm that estimates the principal components of a density matrix using quantum phase estimation and controlled unitary operations. The classical PCA scales poorly with high-dimensional data as it requires eigen-decomposition of the covariance matrix. QPCA can leverage the exponential speedup potential of quantum computing to analyze large datasets. QPCA can be more efficient in processing network traffic data due to the massive feature space. It can preprocess data by reducing irrelevant or redundant features, thereby enhancing efficiency.

### v2

@@ IV-I1 DDoS (Distributed Denial of Service) Attacks
A DDoS attack disrupts the services provided by the systems by flooding them with excessive requests, rendering them unable to function properly. A DDoS attack can have a severe impact on the federated learning process, slowing the model aggregation process or exhausting the limited resources of edge devices [fotse2024federated]. Alhasawi and Alghamdi in [alhasawi2024federated] presented a novel solution, Federated Learning for Decentralized DDoS Attack Detection (FL-DAD) using CNN, to detect DDoS attacks in IoT networks. They also mentioned the challenges while implementing their solutions which include achieving consistent accuracy due to heterogeneous data, resource constraints on edge devices, communication overhead, anomalous data intricacies and computation complexities. In spite of these challenges, FL-DAD constantly achieved an accuracy of 98% for various DDoS classes, outperforming traditional centralized methods.
Similarly, Fotse et al. in [fotse2024federated] developed FedLAD, to detect DDoS attacks in large-scale software-defined networks (SDN). XGBoost classifier was implemented to detect DDoS attacks within the SDN. They implemented three different aggregation strategies and compared the results for three different datasets. The results showed that Astraes and Ranking Client outperform FedAvg in most performance metrics and the results were almost similar to the centralized approach. This method helps in minimal resource consumption by distributing the detection process to multiple controllers enhancing detection speed. A few limitations of FedLAD are model accuracy being dependent on aggregation techniques and the necessity to protect the model against data poisoning threats.
@@ IV-I2 MITM (Man in the Middle) Attacks
MITM attacks are responsible for data theft, unauthorized access and system integrity damage. Various techniques such as ARP spoofing, session hijacking, DNS spoofing and SSL stripping are used for MITM attacks [javeed2020man]. Satpathy et al. in [satpathy2024enhancing] conducted a performance analysis using FL models on Gradient Boosting Machines (FL-GBM) and Long Short-Term Memory (FL-LSTM) combining them with PCA (Principle Component Analysis) for MITM attacks detection. The results showed that FL-GBM outperformed FL-LSTM in terms of accuracy and F1 score.
Similarly, Pathmendre et al. in [pathmendre2024codenexa] presented a novel security framework CodeNexa for MITM attack detection. Existing security mechanisms include hash functions, digital signatures and watermarking to protect the integrity of the data. These traditional methods though effective have several limitations. Hash functions verify static data integrity but are not able to adapt to the dynamic and evolving nature of FL [shamir1979share]. Digital signatures maintain the authenticity of the data but the attacker can intercept and manipulate both data and signature during transmission [rivest1978method]. Similarly, watermarks ensure ownership verification but fail to protect continuous model integrity [samarati2001protecting]. The proposed method does not include cryptography; instead, it enhances model integrity by dynamic metric verification such as accuracy, precision, recall and area under the curve (AUC). During global model aggregation, the stored metrics are compared with newly calculated metrics from local updates before aggregating the model. The model update is accepted only if a certain threshold is met. The metrics need to be transmitted to the global server along with model weights which remains a challenge as it increases the communication overhead.
@@ IV-I3 Botnet Attacks
Alhasawi et al. [alhasawi2024federated] Horizontal Weighted FedAvg Cross-device CICIDS2017 CNN DDoS
Fotse et al. [fotse2024federated] Horizontal Astraes, Ranking Client, FedAVG Cross-silo InSDN, CICDDoS2019, CICDoS2017 XGBoost DDoS
Satpathy et al. [satpathy2024enhancing] Horizontal FedAvg Cross-device UNSW-NB15 GBM, LSTM with PCA MITM
Pathmendre et al. [pathmendre2024codenexa] Horizontal CodeNexa (validation-based) Cross-device MNIST CNN MITM
@@ V Quantum Machine Learning (QML)
Quantum computing is rapidly evolving and getting popular as it leverages the principles of quantum mechanics (i.e., superposition, entanglement, decoherence, and quantum interference) that process information in a different way as compared to classical computers [hdaib2024quantum]. Unlike classical bits (i.e., 0 or 1), quantum bits (qubits) can exist in a superposition of both the states which helps in parallel computation across a vast solution space. A qubit can be in the state $|\psi\rangle=|0\rangle$ or $|\psi\rangle=|1\rangle$, and it can also be in a superposition state, i.e.,
@@
The extreme values of the qubits are 0 and 1, but they can be anything in between, representing a state formed by the combination of 0 and 1. This is a superposition of two values. A qubit is capable of representing a larger amount of data since every state of a qubit can be mapped to different pieces of data. This makes a single qubit solve more computations as compared to many regular classical bits that work simultaneously. Quantum computers rely on quantum Turing machines for computation, in contrast to classical Turing machines in classical computers [deutsch1985quantum]. The runtime of an algorithm is measured in terms of the number of elementary operations $N$ for both classical and quantum computation [montanaro2016quantum]. In classical computing, the same problem can belong to NP but not P, meaning it can be verified in polynomial time but has no known polynomial-time solving algorithm on a classical machine, while it can belong to class BQP in quantum machines when defined on Hilbert spaces [corli2025quantum]. These distinctive features of quantum computing enable them to solve complex optimization and machine-learning problems involving high-dimensional data and intricate patterns.
Quantum computing relies on Dirac notation (also called bra-ket notation) to efficiently describe and manipulate quantum states [dirac1981principles]. A quantum state is written as ket, denoted by $|\psi\rangle$. This represents a column vector in a complex vector space (Hilbert Space). The corresponding bra, denoted by $\langle\phi|$, is a conjugate transpose of a ket. It represents a row vector and is used to form inner products and matrix operations.
@@ V-A Quantum Gates and Circuits
However, quantum gates operate on qubits that can exist in a superposition of 0 and 1. Quantum gates are reversible and linear. The output of these gates gives the probability of qubits being 0 or 1. The gates are classified based on the number of qubits they operate on. Gates that operate on a single qubit are called single-qubit gates, those on two qubits are called two-qubit gates, and those on three qubits are called three-qubit gates [steane1998quantum]. The most known single-qubit gates are Pauli X gate, Pauli Z gate and Hadamard (H) gate:
@@
A quantum circuit is created by combining these gates to get desired outputs. The example of bell state [bell_state] shows how quantum gates can be used to create a quantum circuit and how quantum circuits can be used to perform various operations on quantum bits.
Traditional quantum circuits are static - they are designed to perform fixed computations [10319321]. The concept of parameterized quantum circuits (PQCs), is introduced where certain gates in the circuit contain learnable parameters, similar to weights in classical neural networks. The PQCs are also called variational circuits. In QML, a PQC acts like a model, analogous to a neural network. Classical data are encoded into quantum states using encoding circuits or quantum feature maps [bhowmik2024quantum]. The variational circuit applies transformations based on trainable parameters on these quantum-encoded data. After measurement, classical values are used to calculate a loss function. Various classical optimizers (like gradient descent, Adam) can be used to adjust the parameters based on the loss function. The rotation angles and parameters in the quantum gates in a variational circuit are learned during QML. This learning occurs by minimizing the loss function, similar to classical settings.
@@ V-B Feature Encoding for Quantum Intrusion Detection
The foremost step before implementing QML techniques is to map classical features into quantum states so that QML algorithms can operate on those features. Feature mapping is one of the most crucial steps in QML. Network traffic data used in IDS is high-dimensional, noisy and non-linearly correlated. The quantum feature maps exploit exponentially high-dimensional Hilbert space to encode classical features into quantum states [schuld2019quantum], which allow quantum classifiers or kernels to detect subtle and complex patterns indicative of intrusions or attacks.
@@
The common feature mapping techniques include [ibm_feature]:
@@ V-C QML in Intrusion Detection
With growing interconnectedness between devices [Market], the growth in data has been exponential. In the modern world, we rely on machine learning to help systems make automatic and smart decisions. Machine learning algorithms have made the process of decision-making very fast. One of the most common examples is detecting spam in emails. It could take days for a human to filter spam from a collection of emails by manually checking them whereas a ML algorithm can filter these spam in seconds. But as the amount and complexity of data is constantly growing, classical ML sometimes struggles.
@@
To address this question, it is important to understand the limitations of classical ML algorithms. Classical ML models work by processing combinations of input at a time which makes the model learn patterns point by point [faouzi2023classic]. This becomes computationally heavy if the data is huge and patterns are complex.
@@
These principles led to the integration of quantum algorithms into traditional machine learning workflows [zeguendry2023quantum]. QML models such as quantum support vector machines (QSVMs) [akrom2024quantum], variational quantum circuits (VQCs) [chen2020variational], and quantum kernel methods [schuld2021quantum], can provide exponential or polynomial speed-ups in processing data. In the field of cyber security, QML has been investigated for anomaly detection, traffic pattern analysis and malware detection, where enhanced feature extraction and generalization capabilities are critical [hdaib2024quantum, kukliansky2024network, wang2025network].
As technologies in quantum computing evolve, researchers are exploring the integration of QML into IDS. Most QML-based IDS fall into two main categories - VQC-based and quantum kernel-based [bhowmik2024quantum]. VQC consists of a PQC that uses a sequence of quantum gates controlled by tunable classical parameters. The parameters of PQC are optimized in classical computers to reduce the cost function. On the other hand, kernel-based methods make use of quantum computers to encode the data into high-dimensional Hilbert spaces. The similarities between data points are measured using inner products of quantum states. The quantum kernel captures how similar the two inputs are in this space. A decision boundary is defined by suitable quantum measurement, allowing complex patterns (like anomalies) to be separated linearly in the transformed space.
@@ V-C1 Supervised Quantum Algorithm
  • Quantum Support Vector Machine (QSVM): QSVM is the counterpart of classical SVM. SVM is based on a hyperplane that separates data points of different classes with maximum margin. QSVM is based on a similar approach, but it performs the computation in a high-dimensional Hilbert space using quantum states and quantum kernels. Quantum computers can evaluate the inner products faster, leading to faster evaluation of kernel matrix [rebentrost2014quantum]. The QSVM can encode classical data into exponentially large Hilbert spaces [schuld2019quantum], enabling the separation of classes that are not linearly separable in classical classes. The attacker will make sure that the anomalous data looks like normal data to not get caught, which makes classical SVM separate the normal/anomalous class challenging. However, in QSVM, the encoded data in high-dimensional space makes a clear boundary between normal/anomalous data facilitating more accurate and efficient classification.
@@
  • Quantum Neural Network (QNN): A QNN is a variational circuit with learnable parameters, which are to be optimized while training [qiskit_nn_tutorial]. The gates in variational circuits are controlled by tunable parameters. These tunable parameters are adjusted during training like weights in classical ML. Fig. 8 shows an overview of QML in intrusion detection systems.Designing quantum circuits for intrusion detection include several key steps:
@@
  • Quantum k-Nearest Neighbor (QkNN): The classical kNN algorithm is based on a distance metric (e.g. Euclidean distance). This algorithm identifies the k closest points (neighbors) to a query data point based on a distance metric and uses their label to predict the label of the query data. However, when the data size and dimensionality increase, the classical kNN becomes computationally very expensive due to the need to calculate distances between all the data and a query data point. In QkNN, all training data can be encoded into a superposition of quantum states. Unlike in classical kNN where looping is required on $n$ states one by one, the quantum algorithm can process all of them at once in a single operation using quantum parallelism. The distance in QkNN can be measured using fidelity [basheer2020quantum], which is related to inner products and is given by:
@@
    The fidelity is computed using quantum circuits such as swap tests. Mathematically, fidelity quantifies the overlap between two quantum states, with values ranging from 0 (completely different) to 1 (identical). This measure is crucial for distinguishing subtle differences in network behavior, especially when anomalous traffic closely resembles normal traffic patterns to avoid detection. QkNN can alos be implemented using various distance metrics (such as Hamming and Euclidean distance) [li2022quantum, zardini2024quantum]. In IDS, the query network traffic can be compared with the labeled traffic data using fidelity or various other distance metrics to classify it as normal or anomalous. QkNN can help detect anomalous network data faster as compared to classical kNN due to quantum parallelism.
  • Hybrid Quantum Classical Neural Network (HQCNN): The HQCNN is a combination of quantum and classical neural networks [arthur2022hybrid, ling2024image]. In HQCNN, the feature dimension of classical data is reduced by passing through a classical neural network. The output of the first classical neural network layer acts as input to the quantum circuit. The output is encoded as quantum states using methods like angle or amplitude encoding so that quantum algorithms can operate on it. The quantum operation happens in the second layer. The output of the quantum layer is measured and fed back to the classical neural network layer to get the desired output. This method can efficiently detect anomalies in network traffic where network features have interdependent patterns. Quantum entanglement can help capture dependencies between different types of network behaviors.
@@ V-C2 Unsupervised Quantum Algorithm
  • Quantum Autoencoder (QAE): Classical autoencoders encode classical data into a lower dimension and learn to reconstruct the original classical data from this lower dimension [li2023comprehensive]. The training is done between the input and output of this algorithm to reduce the reconstruction loss. Similarly, QAE encodes the quantum data stored on n qubits into a smaller quantum register of $m<n$ qubits [romero2017quantum](see Fig. 9). In quantum computing, the qubits cannot be simply erased/discarded due to quantum operations being unitary (i.e., reversible). Suppose an initial quantum state living in Hilbert space is given by:
@@
    In [10.1145/1273496.1273497], Aïmeur et al. introduced thee Grover-based quantum subroutines to accelerate classical clustering tasks. These subroutines include quant-find-max (used to find the pair of points in the dataset with maximum distance), quant-find-c-smallest-values (which identifies the $c$ nearest neighbors of a given point), and quant-cluster-median (a novel quantum routine that determines the median of a point set by computing the total distance from each point to all others and then applying Grover’s minimum search to identify the point with the smallest overall distance). Using these primitives, authors developed three clustering algorithms: quantum divisive clustering, $k$-medians and construction of a $c$-neighborhood graph.
@@
    All these algorithms are based on Grover’s algorithm [grover1997quantum] which attains quadratic speedup compared to classical methods. Suppose there are $N$ unsorted items, classically it would take $N/2$ steps to search items on average and $N$ steps in the worst case. Grove’s algorithm can find the solution in $\sqrt{N}$ steps by leveraging quantum parallelism. The ability of quantum computing to handle high-dimensional data and quadratic speedup achieved in Grover’s algorithm shows its potential in network attack detection, particularly suitable for real-time and adaptive security applications.
@@
  • Quantum Ensemble Methods: They extend classical ensemble methods, such as bagging, boosting, and stacking, into the quantum domain [macaluso2020quantum]. In classical learning, many weak learners are used to develop a strong learner that outperforms a single model. Similarly, in quantum computing, quantum ensemble techniques leverage the quantum probabilistic nature and parallelism of quantum computing to produce diverse models and aggregate them to develop a strong model with enhanced performance [schuld2018quantum]. Quantum ensemble methods can capture richer patterns in network traffic using fewer resources compared to the classical model. Classical ensemble methods often struggle with high-dimensional and complex network flow data whereas quantum models can process high-dimensional and complex data in exponentially smaller space which makes quantum ensemble methods more suitable for IDS. It can help reduce false positives and accurately detect anomalies in the network.
@@
  • Quantum Boltzmann Machine (QBM): It is a quantum-enhanced version of the classical Boltzmann machine that leverages the principles of quantum mechanics. The classical energy function is replaced with a quantum Hamiltonian, and the system’s state is described by a thermal quantum density matrix in a QBM [amin2018quantum]. This allows a QBM to model high dimensional probability distribution more efficiently as compared to the classical Boltzmann machine. The patterns in network traffic are very intricate. These patterns can make classical Boltzmann machine hard to learn complex patterns. The QBM can help reduce false positives and detect real-time anomalies due to high-dimensional feature learning which can be very advantageous in IDS.
@@
  • Quantum Generative Adversarial Network (qGAN): The qGAN combines the power of GANs learning paradigm and quantum computation [qiskit_nn_tutorial]. The qGANs use a quantum generator $G_{\Theta}$ to generate synthetic data distributions that closely resemble original data and discriminator $D_{\phi}$ to distinguish between original and generated data distributions. Both generator and discriminator are trained together in a min-max game where $G_{\Theta}$ goal is to fool the discriminator by generating so realistic data that $D_{\phi}$ cannot tell it’s fake and $D_{\phi}$’s goal is to classify the data as real or fake. Lloyd et al. in [lloyd2018quantum] first proposed the concept of qGAN, showing that quantum generators could be more expressive than classical ones, especially when operating in high-dimensional Hilbert spaces. This adversarial learning can be particularly advantageous for intrusion detection, where labeled data is scarce and anomalies are rare, making it suitable for unsupervised or semi-supervised learning. By learning the distribution of normal traffic, qGANs can flag any deviations from normal behavior as a potential threat.
@@
  • Quantum PCA (QPCA): QPCA is a quantum analogue of classical PCA, which is designed to reduce the dimensionality of quantum data while preserving the most significant features. It was first introduced by Lloyd et al. [lloyd2014quantum], who proposed an efficient quantum algorithm that estimates the principal components of a density matrix using quantum phase estimation and controlled unitary operations. The classical PCA scales poorly with high-dimensional data as it requires eigen-decomposition of the covariance matrix. QPCA can leverage the exponential speedup potential of quantum computing to analyze large datasets. QPCA can be more efficient in processing network traffic data due to the massive feature space. It can preprocess data by reducing irrelevant or redundant features, thereby enhancing efficiency.